Calculate the change in the internal energy of a piston expanding against a pressure of 0.75 atm 10.3 L to 22.4 L. In the process, 1 030 J of heat is absorbed.







Consider a 10.0-g CO2 as an ideal gas, expanding isothermally and reversibly from a volume of 4 L to 12 L at 303 K. What is the work done by CO2? Calculate q and ∆E.








Samples of aluminum and of iron, both 1 kg in mass, absorb 9.5 kJ. Which substance would yield a larger change in temperature? Consider the specific heat capacities of aluminum (0.89 J/g-K) and iron (0.45 J/g-K).







1
Expert's answer
2022-03-25T15:37:06-0400

(1) According to first law of thermodynamics

du = dq + dw;

where du = change in internal energy, dq = heat change ( +ve when system gain heat and -ve when system loose heat) , dw = work done by the system(-ve) or work done on the system (+ve)

∆U = q - pdV = 1030 J - 0.75 atm × (22.4 L - 10.3 L) = 1030 J - 7.725 atm L

[∴1Latm=101.33J]

= 1030 J - 7.725 × 101.33 J = 247.2 J

(2) n(CO2) = m(CO2)/Mr(CO2) = 10.0 g / 44.01 g/mol = 0.22722 mol

w = nRTln(V2/V1) = (0.22722 mol) × (8.3145 J/mol K) × 303 K × ln(12 L /4 L) = 629 J

Since the gas expands isothermally, the heat added to the system is

q = w = 629 J

Since q = w in an isothermal process, ∆E = 0

(3) q = m × c × ∆T

9500 J = 1000 g × 0.89 J/g K × ∆T

∆T = 10.67 K (Aluminum)

9500 J = 1000 g × 0.45 J/g K × ∆T

∆T = 21.11 K (Iron)

Iron would yield a larger change in temperature